OpenStudy (anonymous):

plywood glue is stored in a tank has he shape shown. How many cubic feet of glue will this tank hold?

OpenStudy (anonymous):

question is not clear

OpenStudy (anonymous):

what is the shape?

OpenStudy (anonymous):

... shape shown? The volume depends entirely on the shape. If it's rectangular then the volume is simply: \[V = lwh\] (in cubic feet, where l = length, w = width, h = height). If it's cubic and all sides are the same then: \[V = l{^3}\] (where l is any side length). If it's a round take, then it's the product of the area of the bottom and the tank's height, or: \[V = \pi r{^2}h\]

OpenStudy (anonymous):

Sorry I am going to draw right now

OpenStudy (anonymous):



OpenStudy (anonymous):

Ah I see. So you need to combine the volume of the cylindrical part of the tank with the volume of the cone on the top. Volume of cylinder: \[V = \pi r{^2}h\]Volume of cone: \[V = (1/3) \pi r{^2}h\] So... \[V = \pi4{^2}6 + (1/3)\pi 4{^2}4\]\[=(\pi \times 16 \times 6) + ((1/3)\times \pi \times 16 \times 4)\]\[=96\pi + (1/3)64\pi\] Make sense?

OpenStudy (anonymous):

got it Thanks